Shift report 20000904 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Sep 04 2000 - 16:11:40 EDT

  • Next message: http account: "Shift report 20000905 16:00-24:00"

    ------------------------------------------------------------
    Date: Mon Sep  4 16:11:40 2000, Shift: 08:00-16:00
    Supervisor: fv, with: 
    Log book#5 pages 93 - 96
    
    Shift Summary:
    ------------------------------------------------------------
    RHIC up 8:00-12:00
    Some problems re-injection and ramping. Good store
    started a little after 3.00
    
    Beam 4+1 hour during this shift
    
    Continued taken data
    2531  finished 9:30 211k trigger 1
    2532           9:50  91k             (T3 off)
    2533           11:50 158K            (end of store)
    
    T3 had tripped- observed during run 2532.
    
    2534            15:30 69K            (again T3 was not completely on)
    2535 started 15:35
    
    The polarity of D5 was changed to A 2kG (350A) before run 2534.
    
    Run Plan until end of run Tuesday morning (~16 hours left)
    -- MRS 90 deg A polarity. Got about 1.8 at 90B since 40 deg
    measurements.
    -- MRS about 100K zero field
    -- FS  zeror field runs (D1,D2 alternate off) 1/5 field
    -- otherwise continue with FS a 4 deg A pol.
    
    Talked to both Roser & Osaki and agreed since the 
    minimum program presented at Wed. meeting was completed that
    HI run could end at 9/5 despite the problems Thursday 
    and Friday.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Sep 04 2000 - 16:12:08 EDT